Journalist Report – May 23rd

MDRS Crew 195
23 May, 2018

Hump day. The crew continue to evolve as a group. Personality traits surface, morals are tested and maybe even questioned. Debates are heated. It is deliberate of course. We ponder the type of people that may make up the future Mars populations. Would any of us make the cut? Would any of you want to?

The previous few updates have likely not made much sense. I am not suffering from space madness. Big Brother is a brutal censor. Enough said.

The day is hot and clear. The EVAs, as fun and varied as they are, feel like a tease. We whiz by so much photogenic landscape. It’s too much to take in at once. The experience seems a little too fleeting.

One handed ATV driving for a selfie is clearly a no-no and there is the overriding need to return to Hab for repress before our “consumables” run out. Like that annoyingly essential thing called oxygen.

We were given a unique perspective on Mars weather today. Sometimes having limited control over your situation is liberating. Other times it unmasks some cracks. The radio etiquette has shifted from attempted professionalism to requests for urgent ice water enemas on arrival back at Hab. Yep, we’re taking our Sims seriously.

Sol Summary – May 23rd

Crew 195 Daily Summary Report 23.05.2018

MDRS Daily Summary Report for sol 3

Summary Title

Sunny and Hot- cooking on Mars.

Mission Status:


Sol Activity Summary:

Second EVA to Gateway of Lith. Successful sim of medical resuscitation

Look Ahead Plan: EVA to west ridge tomorrow.

Anomalies in work:

ATV 300- continued issues with battery. Unable to hold charge.

Weather:Warmer today. Simulated heat illness

Crew Physical Status:Doing well.

EVA: Tomorrow

Reports to be file:

Operations Report

Photo of the day

Support Requested:


Operations Report – May 23rd

Crew 195 Operations Report 23.05.18

SOL: 3

Name of person filing report: Chris Davis

Non-nominal systems: Continuing run Diesel generator at night. No service rep for Solar yet. Septic serviced yesterday

Notes on non-nominal systems:

Generator (hours run): Overnight approximately 10 plus 4 hours today per Shannon

Solar— SOC 10%

Diesel – Over 50%

Propane – 51 PSI

ATV Fuel (Ethanol free gasoline) – 4 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 90

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1351235

Toilet tank emptied: No
Date last emptied: 5.22.18

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda and 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 1 gallon

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: Honda and ATV 300 out of service.

Deimos rover used: yes
Hours: N/A
Beginning charge: 100%
Ending charge: 69 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes

Beginning charge: 100 %
Ending charge: 71%
Currently charging: Yes

Opportunity rover used: Yes
Beginning charge: 96%
Ending charge: 64%
Currently charging: NA

Curiosity rover used: No
Beginning charge: 100 %
Ending charge: 71%
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Working well

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

EVA Report – May 23rd



Purpose of EVA:

1. Simulation of medical contingency scenarios during EVA. After scouting Gateway to Lith we will return for detailed mapping and navigation of the area.


Successful mapping of Gateway to Lith area.

Destination: Gateway to Lith


Easting: 518300

Northing: 4255600

Participants: Scott D., Mason, Christian, Rae, Ruslana, Chris Davis and Madison Ridgeway

Participants remaining in Hab: Dana Levin, Barry

Road(s) and routes per MDRS Map: Cowdung Road

Mode of travel: ATVs x 3, EVs x3

Copyright © The Mars Society. All rights reserved. | Main Site